# Anonymization Prompt Learning for Facial Privacy-Preserving Text-to-Image Generation

### v2

@@ 1 Introduction
The introduction of text-to-image diffusion models [1, 2, 3] marks a significant milestone in the field of image generation. These models enable the creation of images known for their high resolution, realism, and close alignment with text descriptions. This progress sets new standards for artificial content, and has led to online text-to-image services [2] providing free access to the general public.
@@
However, the ability to forge realistic content with such high degrees of freedom has raised wide social concerns. Among the various problems discussed, the generation of identifiable facial images [4, 5, 6] draws particular attention, as such realistic deepfake generation could lead to violation of privacy. Multiple incidents have occurred recently where text-to-image models were used to create fake images of public figures, which were often indistinguishable from real photos without careful scrutiny. It is anticipated that future iterations of text-to-image models will only intensify the issue, empowering users to realistically create any individual in any setting solely based on textual descriptions.
@@
In wake of this problem, online service providers develop content moderation systems to avoid undesired face generation. For example, OpenAI’s online image generation system DALLE-3 [2] employs a large language model to reject the generation of specific individuals. While effective, such measures introduce extra steps in processing, raising the cost of maintaining a service. To avoid such complicated moderation schemes, researchers also probe into the possibility of removing undesired content from generative models with concept editing [7, 8, 9]. These methods aim at forcing the model to forget sensitive concepts like personal identities [7, 8]. However, most existing methods are constrained to removing only a handful of concepts without severely damaging generation quality. This inability to process more concepts is potentially due to their modification of all or a large subset of network parameters. Consequently, they are insufficient for defending against deepfake image generation, which involves an unlimited number of facial identities.
@@
In this paper, we propose Anonymization Prompt Learning (APL), a simple approach that addresses the deepfake generation issue. Unlike existing methods, APL achieves this without relying on additional content moderation modules or extensive editing of model parameters. To accomplish this, we introduce a learnable soft prompt prefix that is prepended to any prompt input for text-to-image models. Given an identity-specific prompt, we train the prefix such that it forces the diffusion model into generating facial images with inaccurate identities, as demonstrated in fig. 1. This is achieved by denoising facial images so that the output aligns less with the requested identity, but closely with a prompt that describes the person’s prominent facial attribute information. We further train the prefix to be inactive for prompts that do not specify certain identities by aligning the outputs with and without the prefix, thereby maintaining the model’s generation quality. Collectively, a text-to-image diffusion model equipped with the Anonymization Prompt functions almost identically to its original version, except for autonomously generating anonymized images when prompted to depict specific individuals.
@@
Figure 1: Images generated by Realistic Vision v5.1 [10] when prompted with names of public figures. The base model generates highly convincing photos of given identities (top row), which may lead to malicious use. When inserting our Anonymization Prompt (AP) to the model (bottom row), the accuracies of generated identities are significantly reduced.
@@ 2.1 Text-to-Image Diffusion Models
The field of image generation has taken a leap forward thanks to the introduction of diffusion models [11, 12, 13]. Through iterative denoising steps, these models implement a mapping from random noise to the real image distribution, achieving superior performances in terms of image quality and diversity. More recently, Latent Diffusion Model (LDM)  [1] is proposed to operate the diffusion process in a compressed latent space. This practice reduces computational costs and enables the introduction of conditional inputs, such as text descriptions. Taking inspirations from LDM, current text-to-image diffusion models are celebrated for their ability to produce images of high resolution, realism, and close adherence to text descriptions. These advancements have resulted in huge successes like Stable Diffusion [1, 14], Midjourney [15], and DALLE-3 [2], which offer public access to either model parameters or through open API services. Such capabilities enable users to produce an extensive range of realistic creations tailored to their specific requests.
While diffusion models enable numerous creative applications, their utilization raises wide societal concerns regarding potential malicious uses. Particularly, text-to-image models faithfully following textual inputs can be used for inappropriate generations. Among many concerns, one major issue is the generation of high-quality, identifiable facial images, which leads to concerning privacy issues. Several work [4, 5, 6] confirm the capability of models to generate accurate identities when prompted with specific individuals. We target the face generation capability of these general-purposed text-to-image models, aiming to moderate them towards failure at identity-specific generation, but preserving abilities to generate high-quality non-identity-specific content.
@@ 2.2 Content Moderation for Text-to-Image Generation
Online text-to-image service providers employ additional pre-processing or post-processing modules to reject generations deemed inappropriate, including images depicting specific individuals [2]. Meanwhile, researchers of concept editing methods [8, 7, 9, 16] explore the modification of text-to-image models themselves to prevent the generation of designated concepts. ESD [8] leverages a frozen teacher model to provide unconditional guidance for diffusion models towards forgetting of concepts. Concept Ablation [9] proposes model-based and noise-based ablation that both guide the predictions of diffusion models towards pre-defined anchor concepts. Forget-me-not [7] modifies the cross-attention layers of diffusion models to minimize the attention scores between designated concepts and accurate generations. These methods primarily target a limited number of concepts to remove, and tend to suffer from catastrophic forgetting [16] when the number of target concepts increases. This paper demonstrates that a learnable prompt prefix allows for the anonymization of an unlimited number of facial identities, without significantly impacting the quality of generations.
@@ 2.3 Face Anonymization
Face anonymization [17, 18, 19, 20, 21, 22] strives to edit existing facial images to remove identity information, thereby protecting facial privacy. Earlier methods [17, 18] use blurring or pixelation to completely obscure faces. Recent approaches [19, 20, 21, 22] utilize generative models to replace existing faces with others that have different identities. Deepprivacy [19] and CIAGAN [20] use Generative Adversarial Networks [23] conditioned on face landmarks to guide accurate anonymizations. LDFA [21] leverages the powerful inpainting capability of diffusion models to enhance face anonymization for self-driving applications. A recent work [22] proposes to preserve facial attribute information in anonymization with the guidance of a face attribute detection model.
@@ 3.1 Overview
In this section, we introduce the idea of learning to anonymize for diffusion models, and provide information on how to preserve prompt fidelity and generation quality of diffusion models with APL introduced. A detailed pipeline of APL is presented in fig. 2.
@@ 3.2 Text-to-Image Diffusion Models
Diffusion models [11] implement a mapping from random noise to the image distribution by performing a multi-step image denoising. This is achieved by training the model to perform image denoising at different time steps with different noise levels:
$L_{dm}=\mathbb{E}_{x_{t},c,t\sim[1,T]}[||\epsilon_{t}-\epsilon_{\theta}(x_{t},\tau(c),t)||^{2}],$ (1)
where $\epsilon_{t}$ is a sampled Gaussian noise and $t$ is the denoising time step. In the context of Latent Diffusion Models [1], $x_{t}$ is obtained by compressing real images into a latent space and subsequently adding noises. For text-to-image models, $c$ denotes the textual description guiding the image generation, and $\tau$ is implemented by a text encoder.
@@ 3.3 Prompt Learning for Diffusion Models
$[AP]=[V_{1}][V_{2}]...[V_{m}].$ (2)
$[V_{i}]$ is a learnable vector for $i=1,2,\cdots m$, where $m$ is set as a constant. This prefix is prepended to any prompts that enter the diffusion model, which changes the text encoding function into
@@ 3.4 Learning to Anonymize
In existing moderation methods for diffusion models [8, 9, 16], knowledge concerning target concepts are completely removed. However, for the case of face anonymization in text-to-image models, we endeavor to strike a balance between producing completely irrelevant images and retaining original facial attributes, following recent works in face anonymization [22]. In other words, when users request the images of a specific individual, we expect the model to return natural facial images with inaccurate identities, but share similar attributes to the requested identity. This can be especially valuable, as completely altering a person’s appearance could introduce bias, misrepresentation, and degraded user experience.
@@
For example, for $c_{1}$ = "Albert Einstein", we have $c_{2}$ = "A male German-born physicist, bushy mustache, unkempt hair, deep-set eyes. " We note that prompting the model with $c_{2}$ alone almost never leads to the identity specified by $c_{1}$. Therefore, guiding the model to generate images of $c_{2}$ removes the identity, but lead to someone with similar facial attributes that fulfills the description of $c_{2}$. Concretely, we aim at aligning the conditional probabilities $P(x|c_{1})$ and $P(x|c_{2},\rm{not}\ c_{1})$. Following existing work in compositional generalization [24, 25], the latter can be approximated with
@@ 3.5 Quality Preservation
We reveal that given the formulation of $L_{ID}$ above, it is very simple to regularize Anonymization Prompt Learning for above purposes. Specifically, we adopt the LAION-2B [26] dataset to curate a regularization dataset $S_{Reg}=\{(x^{i},c_{1}^{i},c_{2}^{i})\}_{i=1}^{N}$ , which shares the data structure as $S_{ID}$. For each text-image pair $(c,x)$ in LAION, we curate a sample for $S_{Reg}$ by setting $x^{i}=x$ and $c_{1}=c_{2}=c$. We train on these samples with a similar loss function introduced above:
$L_{Reg}=E_{(x,c_{1},c_{2})\sim S_{Reg},t}[||\epsilon_{\theta}(x,\tau_{AP}(c_{1}),t)-\hat{\epsilon}_{\theta}(x,\tau(c_{1}),t)||_{2}^{2}].$ (7)
@@
During training, we mix the samples of $S_{ID}$ and $S_{Reg}$ , and train the model with $L_{APL}=L_{ID}+L_{Reg}$. The pipeline of this process is presented in  fig. 2.
@@ 4.1 Setup
Dataset. We curate dataset $S_{ID}=\{(x^{i},c_{1}^{i},c_{2}^{i})\}_{i=1}^{N}$ for training our Anonymization Prompt. We first collect names of identities that encompass different genders, races, and occupations. With these names as prompts, training images $x^{i}$s are generated with Stable Diffusion v1.5 [1]. We use ChatGPT [27] to generate $c_{2}$ prompts by filling in names and requesting descriptions of the individual’s demographic characteristics and prominent facial attributes. While these responses potentially include inaccurate descriptions, we find this experimental collection procedure to be quite effective in our experiments. We additionally collect another dataset of identities $S_{test}$ for evaluating anonymization on unseen identities. Both $S_{ID}$ and $S_{test}$ contain 100 non-overlapping identities of public figures. Samples of these datasets can be found in the Appendix.
Training. The Anonymization Prompt is trained on the base model of Stable Diffusion v1.5 [1] (SD-v1.5), an advanced text-to-image diffusion model. Hyperparameters are set as $m=10$ and $\alpha=1$. During training, we set the learning rate at $31e-$ 3, and train for $20,000$ steps with batch sizes of 1. We keep all model parameters frozen and only train the learnable prompt.
Evaluating Anonymization. We measure the identity accuracy (ID-ACC) of generated images by calculating the cosine similarity between its Arcface [28] face recognition features and that of the ground truth identity. Lower identity accuracies imply better anonymization. Besides evaluating on SD-v1.5, we demonstrate the transferability of our method with three other text-to-image models, namely SD-v1.4 [1], RealisticVision v5.1 [10], and Stable Diffusion XL (SDXL) [14]. Note that SDXL [14] utilizes a text encoder with different sizes of text embeddings, which bars a direct transfer of our soft prompt. We resolve this issue with a zero-shot prompt transfer method proposed in [29], mapping our prompt to a different embedding space, so that our method remains effective with different model architectures.
Evaluating Quality. We assess the generation quality of both identity-specific and non-identity-specific images. For the former, we calculate FID [30] and facial attribute similarities between faces generated by our model and the original model. For the latter, we generate images with captions from the MS COCO 2014 evaluation set [31], following common practices in generative modeling research [1, 3, 2]. We report CLIP [32] and FID [30] scores to respectively evaluate semantic and quality preservation. Given that our method modifies an existing generative model, we calculate FID with reference images generated by the original model in all cases.
@@ 4.2 Anonymization across Models
SD-v1.5 [1] 0.31566 / 0.37958 0.30254 / 0.36691
SD-v1.5 [1] + APL 0.03924 / 0.10861 0.03970 / 0.10773
SD-v1.4 [1] 0.29911 / 0.36713 0.28837 / 0.35720
SD-v1.4 [1] + APL 0.03893 / 0.10632 0.04647 / 0.11368
RV-v5.1 [10] 0.32205 / 0.37122 0.32071 / 0.36492
RV-v5.1 [10] + APL 0.09215 / 0.14391 0.09033 / 0.14496
SDXL [14] 0.27917 / 0.33371 0.27218 / 0.32830
SDXL [14] + APL 0.07267 / 0.14245 0.07852 / 0.14368
Figure 3: Images generated by three different models when employing the same Anonymization Prompts (AP). Models are prompted to generate Albert Einstein and Taylor Swift, which respectively belong to the training set and the testing set. Note that the AP is only trained on SD-v1.5 [1]. The visual effect demonstrate successful anonymization for any identities across different text-to-image models, consistent with our quantitative results.
Following the training of the Anonymization Prompt, we prepend the learned prefix to different text-to-image diffusion models to evaluate its anonymization performances. We prompt the models to generate images of public figures and calculate identity accuracies (ID-ACC). In this experiment, we use the prompt template "A close-up portrait of [Name]" to ensure that the models generate clear depictions of faces. For each identity, we generate a group of images and calculate both mean and maximum ID-ACCs within the group. Subsequently, we aggregate these scores across all identities to present overall mean and maximum identity accuracies. Results are presented in table 1.
@@
Visually, faces generated by all models are significantly less recognizable, as demonstrated in  fig. 3, consistent with our quantitative results. We thereby confirm the validity of the Anonymization Prompt on four different models, including the state-of-the-art SDXL [14]. This showcases a remarkable transferability of our method, given the prompt is only trained on SD-v1.5.

### v3

@@ 1 Introduction
The introduction of text-to-image diffusion models [ldm, dalle3, imagen] marks a significant milestone in the field of image generation. These models enable the creation of images known for their high resolution, realism, and close alignment with text descriptions. This progress sets new standards for artificial content, and has led to online text-to-image services [dalle3] providing free access to the general public.
@@
However, the ability to forge realistic content with such high degrees of freedom has raised wide social concerns. Among the various problems discussed, the generation of identifiable facial images [antidreambooth, hyperdreambooth, imma] draws particular attention, as such realistic deepfake generation could lead to violation of privacy. Multiple incidents have occurred recently where text-to-image models were used to create fake images of public figures, which were often indistinguishable from real photos without careful scrutiny. It is anticipated that future iterations of text-to-image models will only intensify the issue, empowering users to realistically create any individual in any setting solely based on textual descriptions.
@@
In wake of this problem, online service providers develop content moderation systems to avoid undesired face generation. For example, OpenAI’s online image generation system DALLE-3 [dalle3] employs a large language model to reject the generation of specific individuals. While effective, such measures introduce extra steps in processing, raising the cost of maintaining a service. To avoid such complicated moderation schemes, researchers also probe into the possibility of removing undesired content from generative models with concept editing [forgetmenot, erase, ablating]. These methods aim at forcing the model to forget sensitive concepts like personal identities [forgetmenot, erase]. However, most existing methods are constrained to removing only a handful of concepts without severely damaging generation quality. This inability to process more concepts is potentially due to their modification of all or a large subset of network parameters. Consequently, they are insufficient for defending against deepfake image generation, which involves an unlimited number of facial identities.
@@
In this paper, we propose Anonymization Prompt Learning (APL), a simple approach that addresses the deepfake generation issue. Unlike existing methods, APL achieves this without relying on additional content moderation modules or extensive editing of model parameters. To accomplish this, we introduce a learnable soft prompt prefix that is prepended to any prompt input for text-to-image models. Given an identity-specific prompt, we train the prefix such that it forces the diffusion model into generating facial images with inaccurate identities, as demonstrated in fig.˜1. This is achieved by denoising facial images so that the output aligns less with the requested identity, but closely with a prompt that describes the person’s prominent facial attribute information. We further train the prefix to be inactive for prompts that do not specify certain identities by aligning the outputs with and without the prefix, thereby maintaining the model’s generation quality. Collectively, a text-to-image diffusion model equipped with the Anonymization Prompt functions almost identically to its original version, except for autonomously generating anonymized images when prompted to depict specific individuals.
@@
Figure 1: Images generated by Realistic Vision v5.1 [realisticvision] when prompted with names of public figures. The base model generates highly convincing photos of given identities (top row), which may lead to malicious use. When inserting our Anonymization Prompt (AP) to the model (bottom row), the accuracies of generated identities are significantly reduced.
@@ 2.1 Text-to-Image Diffusion Models
The field of image generation has taken a leap forward thanks to the introduction of diffusion models [ddpm, ddim, adm]. Through iterative denoising steps, these models implement a mapping from random noise to the real image distribution, achieving superior performances in terms of image quality and diversity. More recently, Latent Diffusion Model (LDM)  [ldm] is proposed to operate the diffusion process in a compressed latent space. This practice reduces computational costs and enables the introduction of conditional inputs, such as text descriptions. Taking inspirations from LDM, current text-to-image diffusion models are celebrated for their ability to produce images of high resolution, realism, and close adherence to text descriptions. These advancements have resulted in huge successes like Stable Diffusion [ldm, sdxl], Midjourney [midjourney], and DALLE-3 [dalle3], which offer public access to either model parameters or through open API services. Such capabilities enable users to produce an extensive range of realistic creations tailored to their specific requests.
While diffusion models enable numerous creative applications, their utilization raises wide societal concerns regarding potential malicious uses. Particularly, text-to-image models faithfully following textual inputs can be used for inappropriate generations. Among many concerns, one major issue is the generation of high-quality, identifiable facial images, which leads to concerning privacy issues. Several work [antidreambooth, hyperdreambooth, imma] confirm the capability of models to generate accurate identities when prompted with specific individuals. We target the face generation capability of these general-purposed text-to-image models, aiming to moderate them towards failure at identity-specific generation, but preserving abilities to generate high-quality non-identity-specific content.
@@ 2.2 Content Moderation for Text-to-Image Generation
Online text-to-image service providers employ additional pre-processing or post-processing modules to reject generations deemed inappropriate, including images depicting specific individuals [dalle3]. Meanwhile, researchers of concept editing methods [erase, forgetmenot, ablating, uce] explore the modification of text-to-image models themselves to prevent the generation of designated concepts. ESD [erase] leverages a frozen teacher model to provide unconditional guidance for diffusion models towards forgetting of concepts. Concept Ablation [ablating] proposes model-based and noise-based ablation that both guide the predictions of diffusion models towards pre-defined anchor concepts. Forget-me-not [forgetmenot] modifies the cross-attention layers of diffusion models to minimize the attention scores between designated concepts and accurate generations. These methods primarily target a limited number of concepts to remove, and tend to suffer from catastrophic forgetting [uce] when the number of target concepts increases. This paper demonstrates that a learnable prompt prefix allows for the anonymization of an unlimited number of facial identities, without significantly impacting the quality of generations.
@@ 2.3 Face Anonymization
Face anonymization [blur, pixel, deepprivacy, ciagan, ldfa, attranon] strives to edit existing facial images to remove identity information, thereby protecting facial privacy. Earlier methods [blur, pixel] use blurring or pixelation to completely obscure faces. Recent approaches [deepprivacy, ciagan, ldfa, attranon] utilize generative models to replace existing faces with others that have different identities. Deepprivacy [deepprivacy] and CIAGAN [ciagan] use Generative Adversarial Networks [gan] conditioned on face landmarks to guide accurate anonymizations. LDFA [ldfa] leverages the powerful inpainting capability of diffusion models to enhance face anonymization for self-driving applications. A recent work [attranon] proposes to preserve facial attribute information in anonymization with the guidance of a face attribute detection model.
@@ 3.1 Overview
In this section, we introduce the idea of learning to anonymize for diffusion models, and provide information on how to preserve prompt fidelity and generation quality of diffusion models with APL introduced. A detailed pipeline of APL is presented in fig.˜2.
@@ 3.2 Text-to-Image Diffusion Models
Diffusion models [ddpm] implement a mapping from random noise to the image distribution by performing a multi-step image denoising. This is achieved by training the model to perform image denoising at different time steps with different noise levels:
[…]
where $\epsilon_{t}$ is a sampled Gaussian noise and $t$ is the denoising time step. In the context of Latent Diffusion Models [ldm], $x_{t}$ is obtained by compressing real images into a latent space and subsequently adding noises. For text-to-image models, $c$ denotes the textual description guiding the image generation, and $\tau$ is implemented by a text encoder.
@@ 3.3 Prompt Learning for Diffusion Models
$AP]=[V_{1}][V_{2}]...[V_{m}].$ (2)
$V_{i}]$ is a learnable vector for $i=1,2,\cdots m$, where $m$ is set as a constant. This prefix is prepended to any prompts that enter the diffusion model, which changes the text encoding function into
@@ 3.4 Learning to Anonymize
In existing moderation methods for diffusion models [erase, ablating, uce], knowledge concerning target concepts are completely removed. However, for the case of face anonymization in text-to-image models, we endeavor to strike a balance between producing completely irrelevant images and retaining original facial attributes, following recent works in face anonymization [attranon]. In other words, when users request the images of a specific individual, we expect the model to return natural facial images with inaccurate identities, but share similar attributes to the requested identity. This can be especially valuable, as completely altering a person’s appearance could introduce bias, misrepresentation, and degraded user experience.
@@
For example, for $c_{1}$ = "Albert Einstein", we have $c_{2}$ = "A male German-born physicist, bushy mustache, unkempt hair, deep-set eyes. " We note that prompting the model with $c_{2}$ alone almost never leads to the identity specified by $c_{1}$. Therefore, guiding the model to generate images of $c_{2}$ removes the identity, but lead to someone with similar facial attributes that fulfills the description of $c_{2}$. Concretely, we aim at aligning the conditional probabilities $P(x|c_{1})$ and $P(x|c_{2},\rm{not}\ c_{1})$. Following existing work in compositional generalization [compositional_energy, compositional_diffusion], the latter can be approximated with
@@ 3.5 Quality Preservation
We reveal that given the formulation of $L_{ID}$ above, it is very simple to regularize Anonymization Prompt Learning for above purposes. Specifically, we adopt the LAION-2B [laion] dataset to curate a regularization dataset $S_{Reg}=\{(x^{i},c_{1}^{i},c_{2}^{i})\}_{i=1}^{N}$ , which shares the data structure as $S_{ID}$. For each text-image pair $(c,x)$ in LAION, we curate a sample for $S_{Reg}$ by setting $x^{i}=x$ and $c_{1}=c_{2}=c$. We train on these samples with a similar loss function introduced above:
[…]
@@
During training, we mix the samples of $S_{ID}$ and $S_{Reg}$ , and train the model with $L_{APL}=L_{ID}+L_{Reg}$. The pipeline of this process is presented in  fig.˜2.
@@ 4.1 Setup
Dataset. We curate dataset $S_{ID}=\{(x^{i},c_{1}^{i},c_{2}^{i})\}_{i=1}^{N}$ for training our Anonymization Prompt. We first collect names of identities that encompass different genders, races, and occupations. With these names as prompts, training images $x^{i}$s are generated with Stable Diffusion v1.5 [ldm]. We use ChatGPT [chatgpt] to generate $c_{2}$ prompts by filling in names and requesting descriptions of the individual’s demographic characteristics and prominent facial attributes. While these responses potentially include inaccurate descriptions, we find this experimental collection procedure to be quite effective in our experiments. We additionally collect another dataset of identities $S_{test}$ for evaluating anonymization on unseen identities. Both $S_{ID}$ and $S_{test}$ contain 100 non-overlapping identities of public figures. Samples of these datasets can be found in the Appendix.
Training. The Anonymization Prompt is trained on the base model of Stable Diffusion v1.5 [ldm] (SD-v1.5), an advanced text-to-image diffusion model. Hyperparameters are set as $m=10$ and $\alpha=1$. During training, we set the learning rate at 1e31e-3, and train for $20,000$ steps with batch sizes of 1. We keep all model parameters frozen and only train the learnable prompt.
Evaluating Anonymization. We measure the identity accuracy (ID-ACC) of generated images by calculating the cosine similarity between its Arcface [arcface] face recognition features and that of the ground truth identity. Lower identity accuracies imply better anonymization. Besides evaluating on SD-v1.5, we demonstrate the transferability of our method with three other text-to-image models, namely SD-v1.4 [ldm], RealisticVision v5.1 [realisticvision], and Stable Diffusion XL (SDXL) [sdxl]. Note that SDXL [sdxl] utilizes a text encoder with different sizes of text embeddings, which bars a direct transfer of our soft prompt. We resolve this issue with a zero-shot prompt transfer method proposed in [prompt-transfer], mapping our prompt to a different embedding space, so that our method remains effective with different model architectures.
Evaluating Quality. We assess the generation quality of both identity-specific and non-identity-specific images. For the former, we calculate FID [cleanfid] and facial attribute similarities between faces generated by our model and the original model. For the latter, we generate images with captions from the MS COCO 2014 evaluation set [coco], following common practices in generative modeling research [ldm, imagen, dalle3]. We report CLIP [clip] and FID [cleanfid] scores to respectively evaluate semantic and quality preservation. Given that our method modifies an existing generative model, we calculate FID with reference images generated by the original model in all cases.
@@ 4.2 Anonymization across Models
SD-v1.5 [ldm] 0.31566 / 0.37958 0.30254 / 0.36691
SD-v1.5 [ldm] + APL 0.03924 / 0.10861 0.03970 / 0.10773
SD-v1.4 [ldm] 0.29911 / 0.36713 0.28837 / 0.35720
SD-v1.4 [ldm] + APL 0.03893 / 0.10632 0.04647 / 0.11368
RV-v5.1 [realisticvision] 0.32205 / 0.37122 0.32071 / 0.36492
RV-v5.1 [realisticvision] + APL 0.09215 / 0.14391 0.09033 / 0.14496
SDXL [sdxl] 0.27917 / 0.33371 0.27218 / 0.32830
SDXL [sdxl] + APL 0.07267 / 0.14245 0.07852 / 0.14368
Figure 3: Images generated by three different models when employing the same Anonymization Prompts (AP). Models are prompted to generate Albert Einstein and Taylor Swift, which respectively belong to the training set and the testing set. Note that the AP is only trained on SD-v1.5 [ldm]. The visual effect demonstrate successful anonymization for any identities across different text-to-image models, consistent with our quantitative results.
Following the training of the Anonymization Prompt, we prepend the learned prefix to different text-to-image diffusion models to evaluate its anonymization performances. We prompt the models to generate images of public figures and calculate identity accuracies (ID-ACC). In this experiment, we use the prompt template "A close-up portrait of [Name]" to ensure that the models generate clear depictions of faces. For each identity, we generate a group of images and calculate both mean and maximum ID-ACCs within the group. Subsequently, we aggregate these scores across all identities to present overall mean and maximum identity accuracies. Results are presented in table˜1.
@@
Visually, faces generated by all models are significantly less recognizable, as demonstrated in  fig.˜3, consistent with our quantitative results. We thereby confirm the validity of the Anonymization Prompt on four different models, including the state-of-the-art SDXL [sdxl]. This showcases a remarkable transferability of our method, given the prompt is only trained on SD-v1.5.